OpenStudy (anonymous):

Question of the day :) • Sofia joined a golf club two years ago. She pays an annual membership fee of $1,045 and a greens fee of $35 each time she plays a game of golf. The function below can be used to calculate the total yearly golfing fee, f (g), in dollars. f (g) = 1045 + 35g, where g represents the number of times she played golf during the year. Last year she paid $1,710 as a total golfing fee. For how many games did she pay a greens fee?

OpenStudy (anonymous):

\[f(g) = 1045 + 35g\] g = The number of times she played golf during the year. Last year she paid $1,710 as a total golfing fee. Their asking how many games did she pay a greens fee. You will substitute $1,710 for the "g" in f(g). \[f(1,710) = 1045 + 35g\] You will now subtract 1045 from both sides. \[f(1,710) = 1045 + 35g\] - 1045 - 1045 The 1045 cancels out on the right. \[f(1,710) = 35g\] -1045 \[665 = 35g\] Divide both sides by 35. \[665/35 = 35g/35\] 35 cancels out on the right. \[19 = g\] 0r \[g = 19\] Hope that helps! n_n